# The infinitude of square-free palindromes

### v1

@@ Abstract.
The first author was supported by Australian Research Council Discovery Project DP240100186 and an Australian Mathematical Society Lift-off Fellowship.
The second author was supported by Australian Research Council Grants DE220100859 and DP230100534
@@
footnotetext: Corresponding author: Daniel Johnston (daniel.johnston@unsw.edu.au).footnotetext: Affiliation: School of Science, The University of New South Wales Canberra, Australia.footnotetext: Key phrases: Palindromes, square-free numbers, van der Corput method, exponential sums.footnotetext: 2020 Mathematics Subject Classification: 11A63, 11L07 (Primary) 11N25 (Secondary).
@@ 1.1  Overview
Prior to the 21st century, studies of palindromic numbers were primarily confined to the realm of recreational mathematics. However, in recent years there has been an influx of deep analytic and number-theoretic results regarding palindromes. Examples include the work of Cilleruelo, Luca and Baxter on sums of palindromes [12], or the work of Tuxanidy and Panario [32] on almost-prime palindromes. See also  [25, 2, 3, 24, 15, 13, 14, 33] for other modern works on palindromes. Notably, the recent interest in palindromes is in tandem with other studies of the digital properties of numbers. This includes work on primes with missing digits [27, 28, 23], with preassigned digits [7, 8, 30] and reversed primes [4, 16, 5, 17].
@@
always contains a prime for sufficiently large $x$, even under assumption of the Riemann hypothesis (see e.g. [9, §1.2]).
The problem of whether there exists infinitely many square-free palindromes has often been mentioned in the literature (e.g. [3, p. 10], [24, p. 487], [16, p. 7], [11, p. 371]) and appears as part of Problem 58 in Shparlinski’s list of Problems on Exponential and Character Sums [29]. Here, as usual, a square-free number is a positive integer with no repeated prime factors. In this paper we resolve this problem, and in fact give an asymptotic for the number of square-free palindromes satisfying a standard coprimality condition (see Theorem 1.2).
@@
In the context of Theorem 1.1, it is worth noting that Carillo–Santana [10, Theorem 1.3], as well as Chourasiya and the first author [11, Theorem 1.6], recently and independently proved the existence (and an asymptotic) for cube-free palindromes. However, in both works, it is clear that their techniques are not strong enough to detect square-free palindromes. In another vein, Tuxanidy and Panario recently showed that there are infinitely many palindromes with at most $6$ prime factors [32, Theorem 1.4]. Although this latter result is very deep, their methods are again not strong enough to detect square-free palindromes. In particular, they do not provide a way to bound the number of palindromes with a large square divisor.
To prove Theorem 1.1, we thereby introduce a new approach which combines an elementary argument of Cilleruelo, Luca and Shparlinski [13] with the van der Corput method of exponential sums. Throughout our proof of Theorem 1.1, we prove a series of exponential sum and Fourier-analytic estimates, and new bounds for the number of palindromes with a large square divisor. These subsidiary results may be of independent interest, particularly in the use of weighted sieve methods applied to sets containing palindromes (see [32, § 11]).
@@ 1.2  Main notation and the asymptotic result
Similarly, one can show (see [32, Lemma 9.1])
@@
Notably, the expression in (1.5) has the same form as the asymptotic expression for $k$-free palindromes ($k\geq 3$) given in [11, Theorem 1.6]. Here,
$\frac{1}{\zeta(2)}\prod_{p\mid b^{3}-b}\left(1-\frac{1}{p^{2}}\right)^{-1}$
@@ 1.3  Structure of paper
The proof of Proposition 2.2 is relatively simple, being an elaboration of an elementary argument due to Cilleruelo, Luca and Shparlinski [13, Theorem 1] which reduces counting palindromes with a large square divisor to counting integers in short intervals and arithmetic progressions.
@@
In this section we prove Proposition 2.2. The proof is elementary and inspired by that of [13, Theorem 1]. In particular, the simplest case of [13, Theorem 1] gives an upper bound for the number of square palindromes and we adapt their proof to obtain an upper bound for the number of palindromes with a large square divisor.
@@
is less than $k^{\omega(q)}$, where $\omega(q)$ is the number of distinct prime factors of $q$.
@@ 4.1  The $A$ and $B$ processes
The $A$-process we use is a smooth variation of the usual Weyl-van der Corput inequality. The proof is similar to that of the unsmoothed inequality (see e.g. [19, Lemma 2.5]). However, we include all the details here for clarity.
@@
The $B$-process is then given by the following variation of Poisson summation which, for example, may be readily derived from [6, Lemma 2.9].
@@ 4.2  Oscillatory integral estimates
Using the second mean value theorem, we state two well known estimates on oscillatory integrals. Similar results are given in the literature (e.g. [31, Lemmas 4.3 and 4.5]); however they are not in the right form for our purposes so we provide complete details.

### v2

@@ Abstract.
The first author was supported by Australian Research Council Discovery Project DP240100186 and an Australian Mathematical Society Lift-off Fellowship. The second author was supported by Australian Research Council Grants DE220100859 and DP230100534
@@
footnotetext: Corresponding author: Daniel Johnston (daniel.johnston@unsw.edu.au).footnotetext: Affiliation: School of Science, The University of New South Wales Canberra, Australia.footnotetext: Key phrases: Palindromes, square-free numbers, van der Corput method, exponential sums.footnotetext: 2020 Mathematics Subject Classification: 11A63, 11N25 (Primary) 11L07 (Secondary).
@@ 1.1  Overview
Prior to the 21st century, studies of palindromic numbers were primarily confined to the realm of recreational mathematics. However, in recent years there has been an influx of deep analytic and number-theoretic results regarding palindromes. Examples include the work of Cilleruelo, Luca and Baxter on sums of palindromes [13], or the work of Tuxanidy and Panario [35] on almost-prime palindromes. See also  [26, 3, 4, 25, 16, 14, 15, 31, 1, 36] for other modern works on palindromes. Notably, the recent interest in palindromes is in tandem with other studies of the digital properties of numbers. This includes work on primes with missing digits [28, 29, 24], with preassigned digits [8, 9, 33] and reversed primes [5, 17, 6, 18].
@@
always contains a prime for sufficiently large $x$, even under assumption of the Riemann hypothesis (see e.g. [10, §1.2]).
The problem of whether there exists infinitely many square-free palindromes has often been mentioned in the literature (e.g. [4, p. 10], [25, p. 487], [17, p. 7], [12, p. 371]) and appears as part of Problem 58 in Shparlinski’s list of Problems on Exponential and Character Sums [32]. Here, as usual, a square-free number is a positive integer with no repeated prime factors. In this paper we resolve this problem, and in fact give an asymptotic for the number of square-free palindromes satisfying a standard coprimality condition (see Theorem 1.2).
@@
In the context of Theorem 1.1, it is worth noting that Carillo–Santana [11, Theorem 1.3], as well as Chourasiya and the first author [12, Theorem 1.6], recently and independently proved the existence of infinitely many cube-free palindromes. However, in both works, it is clear that their techniques are not strong enough to detect square-free palindromes. In another vein, Tuxanidy and Panario recently showed that there are infinitely many palindromes with at most $6$ prime factors [35, Theorem 1.4]. Although this latter result is very deep, their methods are again not strong enough to detect square-free palindromes. In particular, they do not provide a way to bound the number of palindromes with a large square divisor.
To prove Theorem 1.1, we thereby introduce a new approach which combines an elementary argument of Cilleruelo, Luca and Shparlinski [14] with the van der Corput method of exponential sums. Throughout our proof of Theorem 1.1, we prove a series of exponential sum and Fourier-analytic estimates, and new bounds for the number of palindromes with a large square divisor. These subsidiary results may be of independent interest, particularly in the use of weighted sieve methods applied to sets containing palindromes (see [35, § 11]).
@@ 1.2  Main notation and the asymptotic result
Similarly, one can show (see [35, Lemma 9.1])
@@
Notably, the expression in (1.5) has the same form as the asymptotic expression for $k$-free palindromes ($k\geq 3$) given in [12, Theorem 1.6]. Here,
$\frac{1}{\zeta(2)}\prod_{p\mid b^{3}-b}\left(1-\frac{1}{p^{2}}\right)^{-1}=\frac{6}{\pi^{2}}\prod_{p\mid b^{3}-b}\left(1-\frac{1}{p^{2}}\right)^{-1}$
@@ 1.3  Structure of paper
Acknowledgements
We thank Aleksandr Tuxanidy for the collegiate and insightful discussions. Notably, Aleksandr informed us that he was working independently on the problem of infinitely many square-free palindromes, and intends to release a paper on this topic containing a quantitatively stronger version of Theorem 2.5. We also thank Igor Shparlinski for some minor remarks on the first version of this manuscript.
@@ Acknowledgements
The proof of Proposition 2.2 is relatively simple, being an elaboration of an elementary argument due to Cilleruelo, Luca and Shparlinski [14, Theorem 1] which reduces counting palindromes with a large square divisor to counting integers in short intervals and arithmetic progressions.
@@
In this section we prove Proposition 2.2. The proof is elementary and inspired by that of [14, Theorem 1]. In particular, the simplest case of [14, Theorem 1] gives an upper bound for the number of square palindromes and we adapt their proof to obtain an upper bound for the number of palindromes with a large square divisor.
@@
is at most $k^{\omega(q)}$, where $\omega(q)$ is the number of distinct prime factors of $q$.
@@ 4.1  The $A$ and $B$ processes
The $A$-process we use is a smooth variation of the usual Weyl-van der Corput inequality. The proof is similar to that of the unsmoothed inequality (see e.g. [20, Lemma 2.5]). However, we include all the details here for clarity.
@@
The $B$-process is then given by the following variation of Poisson summation which, for example, may be readily derived from [7, Lemma 2.9].
@@ 4.2  Oscillatory integral estimates
Using the second mean value theorem, we state two well known estimates on oscillatory integrals. Similar results are given in the literature (e.g. [34, Lemmas 4.3 and 4.5]); however they are not in the right form for our purposes so we provide complete details.
@@ 4.3  Complete exponential sum estimates
Remark.
The start of the above proof (up to (5.5)) may also be obtained via a smoothed-variant of the Erdős-Turán inequality. See for example [30, p. 8] for the usual “unsmoothed” version of this inequality.